OpenStudy (anonymous):

What is the slope of the line passing through the points (1, –5) and (4, 1)?

TheSmartOne (thesmartone):

what is the slope formula

OpenStudy (gabebae):

\[\frac{ y2-y1 }{ x2-x1 }\] If he didn't know..

TheSmartOne (thesmartone):

\(\Large \sf \frac{y_2-y_1}{x_2-x_1}\) where the two points are \(\sf \Large (x_1,y_1)~and~(x_2,y_2)\)

OpenStudy (alexandervonhumboldt2):

the slope formula is \[\frac{ y_1-y_2 }{ x_1-x_2 }~~or~~\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (alexandervonhumboldt2):

substitute your points: \[\frac{ -5-1 }{ 1-4 }=\frac{ -6 }{ -3 }=\frac{ 6 }{ 3 }=???\]

OpenStudy (anonymous):

I'm not sure what you mean by that. Do you mean I should be simplifying that?

OpenStudy (alexandervonhumboldt2):

yes, simplify the 6/3 fraction

OpenStudy (anonymous):

Thanks

TheSmartOne (thesmartone):

so... what did you get??